OpenStudy (anonymous):

Please help: A fastball is hit straight up over home plate. The balls height, "h" (in feet), from the ground is modeled by" h(t)=-16t^2+80t+5 where t is measured in seconds. A) What would be the vertex of this function B) What is the practical meaning of the vertex C) Where is the ball after 1 second D) When is the ball 101 feet above the ground E) What is the horizontal intercept?

OpenStudy (ipwnbunnies):

To find the x-value of the vertex of the parabola (which would be the time, t) use this: \[\frac{-b}{2a}\] Oops, where ax^2 + bx + c = 0 is the general form of a quadratic equation

OpenStudy (anonymous):

okay :)

OpenStudy (ipwnbunnies):

Then you plug that x-value into the equation again, that'll give you the y value, which is the maximum height of the baseball.

OpenStudy (anonymous):

-16t^2+80t+5=0

OpenStudy (ipwnbunnies):

o.o Yes, that's the equation. b = 80 and a = -16

OpenStudy (anonymous):

t=14(10±√105) t≈5.06173769149,−0.06173769149

OpenStudy (anonymous):

thats the first part i think lol

OpenStudy (ipwnbunnies):

Idk what you just did. o.o

OpenStudy (anonymous):

lmao.

OpenStudy (ipwnbunnies):

-b/2a finds the x-value for the vertex. We're not using quadratic formula.

OpenStudy (anonymous):

-2.5?